# Reward-Augmented Data Enhances Direct Preference Alignment of LLMs

### v2

@@ A.1 Setup
For hyperparameters, we utilize a KL regularization coefficient of $\beta=0.01$ in DPO, and we adopt the AdamW optimizer (Loshchilov, 2017). The batch size is set to $128$, with a learning rate of $75e-$ 7 and a warmup ratio of $0.1$. Furthermore, we observe that for models such as Qwen2-7B-Instruct and Gemma-2-9B-It on UltraFeedback, as well as Llama-3-8B-Instruct on on-policy data, both DPO and our proposed method yield improved performance when employing the conservative DPO (cDPO) technique (Mitchell, 2023). Consequently, for these models, we set the label smoothing hyperparameter from the Alignment Handbook (Tunstall et al., 2023a) to $0.3$, while keeping it at 00 for the remaining models.
A.2 Full Results

### v3

@@ B.1 Setup
For hyperparameters, we utilize a KL regularization coefficient of $\beta=0.01$ in DPO, and we adopt the AdamW optimizer (Loshchilov, 2017). The batch size is set to $128$, with a learning rate of $75e-$ 7 and a warmup ratio of $0.1$. Furthermore, we observe that for models such as Qwen2-7B-Instruct and Gemma-2-9B-It on UltraFeedback, as well as Llama-3-8B-Instruct on on-policy data, both DPO and our proposed method yield improved performance when employing the conservative DPO (cDPO) technique (Mitchell, 2023). Consequently, for these models, we set the label smoothing hyperparameter from the Alignment Handbook (alignment_handbook2023) to $0.3$, while keeping it at 00 for the remaining models.
B.2 Full Results
@@
We also provide the full comparison results with reward-augmented methods in Table 12.
Zephyr-SFT DPO DPA SteerLM NCA-P NCA-R INCA-P INCA-R Ours
LC Win Rate 6.21 11.60 11.13 - 11.50 12.87 13.68 14.83 16.66
Win Rate 3.94 8.58 10.58 8.21 8.43 9.56 11.00 11.34 13.37
Avg. Len. 893 1240 1671 1585 1287 1364 1449 1338 1812
Table 12: Full comparison results with reward-augmented methods.
B.3 More Ablations
Controllable Generation with Prompt.
$g=10$
$g=8$
$g=6$
UF ($g=10$)
LC WR 31.17 28.66 25.56 24.44
WR 27.58 25.57 18.88 20.75
Table 13: Performance when conditioned on different goal rewards in the inference prompt.
In Table 13, we ablate how generations differ when changing the goal rewards in the system prompt. We observe that the AlpacaEval 2.0 scores of the Qwen2-7B-It+DPO (RA) model change accordingly as $g$ varies. However, using the same $g=10$ prompt during inference for the Qwen2-7B-It+DPO (UF) model fails to give competitive results, indicating that our method is superior not only because of the additional system prompt.
Benefits of Learning from High-Quality Rejected Responses.
Using the UltraFeedback dataset, we construct two reward-augmented preference datasets by filtering out augmented data based on rejected responses with low and high reward values, respectively. Compared to our method, these datasets isolate the impact of excluding low- and high-reward rejected responses as goals. The evaluation results on AlpacaEval 2.0 are presented in Table 14. Learning from rejected high-reward samples demonstrates superior performance compared to the approach that excludes these samples.
Qwen2-7B-It +DPO (UF) +DPO (RA)
+DPO (RA
filter high)
+DPO (RA
filter low)
LC Win Rate 20.93 21.46 31.17 29.36 31.81
Win Rate 18.22 19.35 27.58 27.04 27.28
Table 14: Ablation on the benefits of learning from high-quality rejected responses.
Impact of the Reward Scale.
For the UltraFeedback dataset that contains response rewards in the range of 1-10, we relabel them to be in the range of 1-5 and 1-100 with linear transformation. Our method followed by DPO is then applied on these different scaled datasets. The results are shown in Table 15. It can be observed that our method is robust to the reward scales. Since our main experiments use the default 1-10 scale as in UltraFeedback, it is likely that the performance can be further boosted, e.g., by adopting the 1-100 scale.
Qwen2-7B-It +DPO (UF) +DPO (RA, 5) +DPO (RA 10) +DPO (RA 100)
LC Win Rate 20.93 21.46 29.85 31.17 31.81
Win Rate 18.22 19.35 26.12 27.58 27.96
Table 15: Ablation on the impact of the reward scale demonstrates the robustness of our method.